Find the equation of the line that passes through the point (2,3,4) and is perpendicular to the plane 2x-y + 3z = 4

a. x=4+2t , y=2-t, z=7-3t

b. x=2+2t , y=3-t, z=4+t

c. x=2-2t , y=-3+t, z=4-3t

d. x=-2+4t , y=5-2t, z=-2+6t

e. another solution
Answer #1

we are given plane as

2x-y + 3z = 4

and it is passing through (2,3,4)

now, we can get equation of line

(x,y,z)=(2,3,4)+t(2,-1,3)

(x,y,z)=(2+2t,3-t,4+3t)

so, we get

x=2+2t,y=3-t,z=4+3t............Answer